Please answer questions 1 and 2
on The rate constant for the reaction answered A + B->C out of 1.0 is 4.1 x104 U(mol s) at 593 K. The activation energy is 101 kj/mol. question Determine the temperature when the rate constant is 1.2 x10-3 L/(mol s)- Answer: 2 One (often inaccurate) rule of thumb in biology is that a 10 °C change increases the rate by a factor of ten. If the lower temperature is 3 °C, what activation energy does this correspond? Report your answer in kj/mol. on et answered ed out of 1.0 g question Answer tion 3 The reaction rate for a chemical reaction was measured as a function of temperature. Determine the activation energy for the reaction. Report your answer in k/mol answered ked out of 1.0 Temp/K Rate constant /arb.
Answer #1

According to arrhenius equation

K = Koe-EA/RT

K = rate constant

T = temperature

EA = activation energy

for 2 different temperature ,the equation can be rearranged as:

ln(K2/K1) = EA/R[1/T1 - 1/T2]

ln{1.2 x 10-3/4.1 x 10-4 } = 10100J/mol / 8.314J/mol.K[1/593K - 1/T2]

T2 = 1246.48 K

b)

3oC = 273 + 3 = 276K

10 oC increment = 276 + 10 = 286K

ln(r2/r1) = EA/8.314J/mol.K[1/276K - 1/286K] given r2 = 10 r1

ln10 x 8.314 /[1/276K - 1/286K] = EA

151112.65 J/mol = 151.113 kJ/mol = EA